OpenStudy (fwer74):

Match the rectangles formed by the sets of points to their corresponding areas. Tiles A(-9, 8), B(-5, 5), C(1, 13), D(-3, 16) 50 square units E(30, 20), F(39, 29), G(49, 19), H(40, 10) 300 square units I(-6, 2), J(2, 2), K(2, -8), L(-6, -8) 100 square units M(5, 5), N(11, 5), O(11, -5), P(5, -5) 80 square units Q(10, 0), R(15, 5), S(25, -5), T(20, -10) U(0, 5), V(15, 20), W(25, 10), X(10, -5)

OpenStudy (eliesaab):

AreaOfRectangle[(m,n},(p,q),(r,s),(t,u)] is given by \[ \sqrt{\left| m-p\right| ^2+\left| n-q\right| ^2} \sqrt{\left| p-r\right| ^2+\left| q-s\right| ^2} \]

OpenStudy (eliesaab):

Compute each one of them and match them